# LocoTouch: Learning Dexterous Quadrupedal Transport with Tactile Sensing

### v1

@@ Abstract
Quadrupedal robots have demonstrated remarkable agility and robustness in traversing complex terrains. However, they remain limited in performing object interactions that require sustained contact. In this work, we present LocoTouch, a system that equips quadrupedal robots with tactile sensing to address a challenging task in this category: long-distance transport of unsecured cylindrical objects, which typically requires custom mounting mechanisms to maintain stability. For efficient large-area tactile sensing, we design a high-density distributed tactile sensor array that covers the entire back of the robot. To effectively leverage tactile feedback for locomotion control, we develop a simulation environment with high-fidelity tactile signals, and train tactile-aware transport policies using a two-stage learning pipeline. Furthermore, we design a novel reward function to promote stable, symmetric, and frequency-adaptive locomotion gaits. After training in simulation, LocoTouch transfers zero-shot to the real world, reliably balancing and transporting a wide range of unsecured, cylindrical everyday objects with broadly varying sizes and weights. Thanks to the responsiveness of the tactile sensor and the adaptive gait reward, LocoTouch can robustly balance objects with slippery surfaces over long distances, or even under severe external perturbations.
Figure 1: We present LocoTouch, a system for quadrupedal robots to learn dynamic transport of unsecured objects using tactile sensing. Equipped with a custom distributed tactile sensor on its back, LocoTouch enables the robot to robustly transport a wide range of unsecured everyday objects with diverse physical properties.
Keywords: Legged Robots, Tactile Sensing, Reinforcement Learning
@@ 4 Distributed Tactile Sensing and Simulation
5 Learning Tactile Policy with Adaptive Gait Reward
@@ 5.2 Teacher Policy Training
$r_{\text{alive}}$ is a constant that ensures the overall reward remains positive. $r_{\text{task}}$ encourages the robot to track the velocity command and balance the object. $r_{\text{reg}}$ is a set of regulatory terms for stable and smooth locomotion [3]. To mitigate the sim-to-real gap caused by foot deformation and enable stable deployment in the real world, we define $r_{\text{drag}}$ as a sum of four binary penalty terms, one for each leg. Each penalty activates when a foot is near the ground while exhibiting high lateral velocity (details in Appendix D.2). Finally, $r_{\text{gait}}$ is an adaptive term that promotes symmetric and stable gait patterns. We find this term to be essential for learning a robust and reliable transport policy, and we discuss it in detail in Section 5.3.
@@ 5.3 Adaptive Gait Reward
where the first term encourages synchronized contact states between diagonal legs, and the second term promotes alternating contact states between lateral legs. To further guide the learning of symmetric air times and encourage a relatively low gait frequency, we define an adaptive coefficient $\gamma_{\text{sym}}$ for the diagonal term. We briefly discuss the construction of $\gamma_{\text{sim}}$ here, and include the detailed formulation in Appendix D.2. We define $\gamma_{\text{sim}}$ as:
@@ 5.4 Student Policy Distillation
[…]
@@ 6.3 High-Fidelity Tactile Simulation with Expanded Collision Model
Figure 6: Comparison of tactile maps generated by different simulation methods with real sensor signals. Object poses are ($x=0.01\,\mathrm{m},\ y=0,\ \text{yaw}=0^{\circ}$) and ($x=0,\ y=0,\ \text{yaw}=30^{\circ}$).
@@ 8 Limitations
Appendix A Extended Related Works
Symmetric Locomotion Learning. Symmetry is a key feature of healthy gaits, promoting locomotion stability and adaptability while reducing the risk of falling [54]. Different methods that enforce symmetric gaits for locomotion policies trained using reinforcement learning have been evaluated as effective [55, 56, 57, 58]. Their common idea is to take advantage of the robot’s kinematic symmetry and design mechanisms such as mirror symmetry loss or symmetry-based data augmentation, to encourage policies to exhibit spatial symmetry in the state-action space [55, 57]. However, these approaches face several limitations. First, they are often robot-specific because the assumed kinematic symmetry varies across platforms. For example, ETH Anymal [59] is symmetric both left-right and front-rear, while MIT Cheetah 3 robots [60] and Unitree Go1 are only left-right symmetric. This requires a careful, manual specification of the mirror symmetry function for robots with new kinematics. Second, these methods generally assume that robot dynamics and sensor configurations share the same symmetry as the robot kinematics, which rarely holds in real robots. Third, they aim to enforce symmetry across the entire state-action space, which becomes increasingly challenging in problem settings with high-dimensional states. Lastly, they only encourage gait symmetry in kinematic trajectories, neglecting timing aspects such as balanced swing-stance durations. In contrast, our proposed symmetricity function directly encourages gait-level symmetry, independent of robot morphology, sensor layout, or environmental configuration. This makes it broadly applicable across diverse tasks and platforms.
Appendix B Details for Distributed Tactile Sensing
B.1 Expanded Collision Model Calibration
@@
B.2 Latency Calibration
We use a force gauge to press on our distributed tactile sensor and record the readings from both devices. As shown in Figure 11, the tactile sensor readings exhibit a delay of approximately $0.02,\mathrm{s}$ compared to the force gauge. The force gauge, operating at $80,\mathrm{Hz}$ with negligible latency, serves as a reliable reference. Based on this comparison, we estimate the tactile sensor’s latency to be in the range of 0.020.020.020.020.$0.04,\mathrm{s}$, which is relatively low compared to that of a depth camera. Using a higher-frequency force gauge would allow for more precise latency calibration, and we plan to adopt such equipment in future work.
@@ B.2 Latency Calibration
Appendix C System Setup
Figure 12: System setup in the real world.
Appendix D Details for Tactile Policy Training
@@
Reference Frame Robot frame $(\cdot)^{\text{r}}$
World frame $(\cdot)^{\text{w}}$
@@
Quaternion $\bm{Q}_{\text{o}}^{\text{r}}$
Euler angle $\bm{\theta}_{\text{o}}^{\text{r}}$ $rad$
@@
Quaternion $\bm{Q}^{\text{w}}$
Euler angle $\bm{\theta}^{\text{w}}$ $rad$
@@
Robot Joint Joint position $\bm{q}$ $rad$
@@
Foot contact force $\bm{F}_{\text{f}}^{\text{w}}$ $N$
@@
D.1 Teacher Policy Training
@@
Training Environment. As Figure 13 illustrates, we train the teacher policy in IsaacLab [52] using 4096 parallel simulation environments. The robot includes collision models for both the back shell and the tactile sensors; however, force tracking for the tactile sensors is disabled during this phase, as the teacher policy does not rely on tactile signals. To improve robustness, we apply environment randomization using the terms listed in Table 3.
@@
$yaw\in[-30,30]$
Mass $[0.5,2.5]$ $kg$
Friction Coefficient $[0.3,1.0]$
Robot Trunk Mass $[4.2,6.2]$ $kg$
Trunk Friction Coefficient $[0.3,1.0]$
Foot Friction Coefficient $[0.6,1.5]$
Initial Delta Joint Position $[-0.03,0.03]$ $rad$
Initial Joint Velocity $[-0.1,0.1]$ $rad/s$
@@
Policy Network. The teacher policy (actor) is implemented as a multi-layer perceptron (MLP) that takes as input the state sequence $o_{(t-H)\dots t}\in\mathbb{R}^{340}$. The network consists of four fully connected layers with hidden dimensions of $512$, $256$, and $128$, each followed by an ELU activation. It outputs an action vector $a_{t}\in\mathbb{R}^{12}$, representing the target delta joint positions. This output is clipped to the range $[-100,100]$ for stable training and scaled by a factor $\alpha^{\text{action}}=0.25$, producing the target joint positions as $q_{t}^{\text{target}}=\alpha^{\text{action}}\cdot a_{t}+q^{\text{default}}$, where $q^{\text{default}}$ is the default joint positions of the robot. A PD controller with gains $K_{p}=25$ and $K_{d}=0.5$ is then used to track the target joint positions $q_{t}^{target}$.
Post-Processing for Observations. To improve robustness and narrow the domain gap for the tactile student policy which relies on partial observations, we add noise to the actor’s input during training. Each processed observation term is computed as $o_{i}^{input}=\alpha_{i}^{input}\cdot(o_{i}^{raw}+o_{i}^{noise})$, where $o_{i}^{raw}$ is the $i$-th observation term from simulation, $o_{i}^{noise}$ is the added noise, $\alpha_{i}^{input}$ is a scaling factor, and $o_{i}^{input}$ is the processed observation term fed into the policy. Notably, for the object orientation, noise is first added to the Euler angles before converting them to quaternions which are used as the final input. The details of each observation term, including noise ranges and scaling factors, are summarized in Table 4.
@@
Object State Position $[-0.01,0.01]$ 1.0 $m$
Linear Velocity $[-0.2,0.2]$ 0.5 $m/s$
Orientation $[-0.05,0.05]$ 1.0 $rad$
Angular Velocity $[-0.2,0.2]$ 0.25 $rad/s$
Robot Proprioception Projected Gravity Vector $[-0.05,0.05]$ 1.0 -
Base Angular Velocity $[-0.2,0.2]$ 0.25 $rad/s$
Joint Position $[-0.01,0.01]$ 1.0 $rad$
Joint Position $[-1.5,1.5]$ 0.05 $rad/s$
@@
Last Action - 1.0 $rad$
D.2 Reward Design

### v2

# LocoTouch: Learning Dynamic Quadrupedal Transport with Tactile Sensing

@@ Abstract
Quadrupedal robots have demonstrated remarkable agility and robustness in traversing complex terrains. However, they struggle with dynamic object interactions, where contact must be precisely sensed and controlled. To bridge this gap, we present LocoTouch, a system that equips quadrupedal robots with tactile sensing to address a particularly challenging task in this category: long-distance transport of unsecured cylindrical objects, which typically requires custom mounting or fastening mechanisms to maintain stability. For efficient large-area tactile sensing, we design a high-density distributed tactile sensor that covers the entire back of the robot. To effectively leverage tactile feedback for robot control, we develop a simulation environment with high-fidelity tactile signals, and train tactile-aware transport policies using a two-stage learning pipeline. Furthermore, we design a novel reward function to promote robust, symmetric, and frequency-adaptive locomotion gaits. After training in simulation, LocoTouch transfers zero-shot to the real world, reliably transporting a wide range of unsecured cylindrical objects with diverse sizes, weights, and surface properties. Moreover, it remains robust over long distances, on uneven terrain, and under severe perturbations.
Figure 1: We present LocoTouch, a system for learning tactile-aware quadrupedal policies with zero-shot sim-to-real transfer. Equipped with a custom distributed tactile sensor on the back, LocoTouch enables quadrupedal robots to robustly transport a wide range of unsecured everyday objects with diverse physical properties.
Keywords: Tactile Quadrupedal Policy, Tactile Sim-to-Real, Legged Robots
@@ 4 Distributed Tactile Sensing and Simulation
5 Tactile-Aware Quadrupedal Policy Learning
@@ 5.2 Teacher Policy Training
$r_{\text{alive}}$ is a constant that ensures the overall reward remains positive. $r_{\text{task}}$ encourages the robot to track the velocity command and balance the object. $r_{\text{reg}}$ is a set of regulatory terms for stable and smooth locomotion [3]. To mitigate the sim-to-real gap caused by foot deformation and enable stable deployment in the real world, we define $r_{\text{drag}}$ as a sum of four binary penalty terms, one for each leg. Each penalty activates when a foot is near the ground while exhibiting high lateral velocity (details in Appendix C.4). Finally, $r_{\text{gait}}$ is an adaptive term that promotes symmetric and stable gait patterns. We find this term to be essential for learning a robust and reliable transport policy, and we discuss it in detail in Section 5.3.
@@ 5.3 Adaptive Gait Reward
where the first term encourages synchronized contact states between diagonal legs, and the second term promotes alternating contact states between lateral legs. To further guide the learning of symmetric air times and encourage a relatively low gait frequency, we define an adaptive coefficient $\gamma_{\text{sym}}$ for the diagonal term. We briefly discuss the construction of $\gamma_{\text{sim}}$ here, and include the detailed formulation in Appendix C.4. We define $\gamma_{\text{sim}}$ as:
@@ 5.4 Student Policy Distillation
After training the teacher policy that relies on object state obtained from simulation, we aim to train a student policy based on onboard tactile sensing for deployment. The only input difference between the teacher and student policies is that the student policy observes tactile map while the teacher policy observes object state sequence. As shown in Figure 2, following a similar monolithic architecture as in [2], we use a convnet-GRU encoder to encode the tactile map and feed the tactile embedding, along with other observations, into the MLP-based policy backbone. We adopt a two-stage distillation process [53], first collecting a dataset of 400k state-action pairs from the teacher to pre-train the student policy via Behavior Cloning (BC). We then refine the student policy using DAgger [50], training it over 7 iterations, with each iteration collecting 200k additional state-action pairs. To better match real-world conditions, we also measure the tactile signal latency ($0.02\,\text{--}\,0.04\text{s}$) and simulate this delay during training.
@@ 6.3 High-Fidelity Tactile Simulation with Expanded Collision Model
[…]
@@ 8 Limitations
9 Acknowledgments
The work is partially supported by Google Deepmind with an unrestricted grant.
Appendix A Related Work on Symmetric Locomotion Learning
Symmetry is a fundamental property of healthy gaits, contributing to locomotion stability, adaptability, and reduced risk of falling [54]. Various methods have been proposed to enforce gait symmetry in reinforcement learning-based locomotion policies and have demonstrated notable effectiveness [55, 56, 57, 58]. The common principle behind these approaches is that policies should produce symmetric actions when observing symmetric states, where the definition of action symmetry is derived from the kinematic symmetry of the robot. To achieve this, previous work introduces mechanisms such as mirror symmetry losses [55] and symmetry-based data augmentation [57, 58], enforcing them over the entire state–action space.
Despite their effectiveness, these approaches have several limitations:
  1. 1.
    Robot-specific assumptions. The definition of mirror symmetry depends heavily on robot morphology. For example, ETH Anymal [59] is symmetric both left-right and front-rear, while MIT Cheetah 3 robots [60] and Unitree Go1 are only left-right symmetric. This requires a careful and manual specification of symmetry mappings for each new platform.
  2. 2.
    Mismatch between kinematic and physical symmetry. These methods often assume that robot dynamics and sensor configurations share the same symmetry as kinematics. In practice, real robots rarely satisfy this assumption due to asymmetries in hardware, sensors, or payload.
  3. 3.
    High-dimensional state challenges. Enforcing symmetry over the entire state–action space becomes increasingly difficult as observation spaces grow, particularly for policies that incorporate high-dimensional sensory inputs such as images.
  4. 4.
    Neglect of temporal symmetry. These methods promote symmetry only in kinematic trajectories, overlooking temporal aspects such as balanced swing and stance durations across legs.
Our key insight is that robots should exhibit symmetric gait behaviors when task commands (e.g., velocity commands) remain constant, and that this requirement can be relaxed under degraded task performance. For example, when the robot or its carried object experiences a sudden external disturbance, it may need to adjust step timing asymmetrically to maintain balance. In our formulation, gait symmetry is defined as equal swing or stance durations within each pair of legs (e.g., diagonal pairs in trotting gaits). Based on this insight, we propose a symmetricity function that evaluates symmetry directly in terms of gait behavior (swing–stance phases) rather than kinematic states or actions. This approach is independent of robot morphology, sensor configuration, and environmental conditions, making it broadly applicable across tasks and platforms. Moreover, by integrating this symmetricity function into the reward, we directly encourage gait-level symmetry through policy gradients, without introducing additional loss terms or complicating the learning paradigm.
For clarity, the key comparison metrics with prior methods are summarized in Table 2.
Table 2: Comparison of methods for symmetric locomotion learning.
Prior Works [55, 56, 57, 58] LocoTouch
Symmetry Basis Robot Kinematics Gait Definition
Symmetry Condition States or Observations Commands and Task Performance
Symmetry Target Whole-Body Actions Foot Contact Patterns
Enforcement Auxiliary Loss or Augmentation Reward Function
Appendix B Details for Distributed Tactile Sensors
B.1 Principle and Fabrication
As introduced in Section 4, our tactile sensor leverages the principle of piezoresistive sensing, where the electrical resistance of certain materials changes under mechanical deformation. Similar to tactile gloves [51] and tactile grippers [35], we use Velostat (Linqstat) as the piezoresistive material.
However, our conductive layer design differs significantly from them [51, 35], leading to substantial improvements: (a) High Sensitivity. Previous designs rely on conductive threads, which make only minimal contact with the piezoresistive layer, thereby limiting sensitivity. In contrast, we use a thin Faraday conductive fabric that provides a much larger contact area. This design greatly enhances sensitivity, allowing the sensor to reliably detect even a lightweight object such as a penny without external pressing. (b) Scalable Fabrication. Conductive threads require manual placement, which is time-consuming, imprecise, and prone to air bubbles within the adhesive layers. By contrast, our method employs laser cutting of the conductive fabric, enabling efficient, precise, and highly scalable fabrication simply by stacking pre-cut layers.
Figure 9: Materials, cutting files, and fabrication procedures for building the tactile sensors. The illustrated sensor shape is for demonstration and is differ from the final version used in LocoTouch.
As shown in Figure 9, the main components of the tactile sensor are: (1) Velostat, which serves as the piezoresistive material; (2) Faraday conductive fabric, which interfaces with Velostat; (3) Transfer tape, which fastens and encapsulates the layers.
The fabrication process consists of the following steps:
  1. 1.
    Attach the conductive fabric onto the transfer tape.
  2. 2.
    Laser cut the rows or columns on the conductive fabric, carefully tuning laser power and speed to avoid cutting through the transfer tape. Fiber laser would be ideal since it doesn’t cut through the transfer tape.
  3. 3.
    Laser cut the outer shape of the sensor on both the conductive fabric and the transfer tape.
  4. 4.
    Remove the excess conductive fabric, leaving only the patterned rows or columns adhered to the transfer tape.
  5. 5.
    Attach external cables to one end of the rows and columns for readout.
  6. 6.
    Laser cut the Velostat layer and stack it with the two conductive layers prepared above.
For additional resources on piezoresistive tactile sensors, we refer readers to the open-source tactile-skin toolkit.
B.2 Expanded Collision Model Calibration
@@
B.3 Latency Calibration
[…]
@@ B.3 Latency Calibration
Appendix C Details for Tactile Policy Training
@@
Reference Frame Robot frame $(\cdot)^{\text{r}}$ -
World frame $(\cdot)^{\text{w}}$ -
@@
Quaternion $\bm{Q}_{\text{o}}^{\text{r}}$ -
Euler angle $\bm{\theta}_{\text{o}}^{\text{r}}$ radrad
@@
Quaternion $\bm{Q}^{\text{w}}$ -
Euler angle $\bm{\theta}^{\text{w}}$ radrad
@@
Robot Joint Joint position $\bm{q}$ radrad
@@
Foot contact force $\bm{F}_{\text{f}}^{\text{w}}$ NN
@@
C.1 Basic Configurations in Simulation
@@
Training Environment. As Figure 13 illustrates, we train the teacher policy in IsaacLab [52] using 4096 parallel simulation environments. The robot includes collision models for both the back shell and the tactile sensors; however, force tracking for the tactile sensors is disabled during this phase to speed up the contact computation, as the teacher policy does not rely on tactile signals. To improve robustness, we apply domain randomization to the environment, using the terms listed in Table 4.
@@
$yaw\in[-30,30]$ °
Mass $[0.5,2.5]$ kgkg
Friction Coefficient $[0.3,1.0]$ -
Robot Trunk Mass $[4.2,6.2]$ kgkg
Trunk Friction Coefficient $[0.3,1.0]$ -
Foot Friction Coefficient $[0.6,1.5]$ -
Initial Delta Joint Position [0.03,0.03][-0.03,0.03] radrad
Initial Joint Velocity [0.1,0.1][-0.1,0.1] $rad/s$
@@
Policy Network. The teacher policy (actor) is implemented as a multi-layer perceptron (MLP) that takes as input the state sequence $o_{(t-H)\dots t}\in\mathbb{R}^{340}$. The network consists of four fully connected layers with hidden dimensions of $512$, $256$, and $128$, each followed by an ELU activation. It outputs an action vector $a_{t}\in\mathbb{R}^{12}$, representing the target delta joint positions. This output is clipped to the range [100,100][-100,100] for stable training and scaled by a factor $\alpha^{\text{action}}=0.25$, producing the target joint positions as $q_{t}^{\text{target}}=\alpha^{\text{action}}\cdot a_{t}+q^{\text{default}}$, where $q^{\text{default}}$ is the default joint positions of the robot. A PD controller with gains $K_{p}=25$ and $K_{d}=0.5$ is then used to track the target joint positions $q_{t}^{target}$.
Post-Processing for Observations. To improve robustness and narrow the domain gap for the tactile student policy which relies on partial observations, we add noise to the actor’s input during training. Each processed observation term is computed as $o_{i}^{input}=\alpha_{i}^{input}\cdot(o_{i}^{raw}+o_{i}^{noise})$, where $o_{i}^{raw}$ is the $i$-th observation term from simulation, $o_{i}^{noise}$ is the added noise, $\alpha_{i}^{input}$ is a scaling factor, and $o_{i}^{input}$ is the processed observation term fed into the policy. Notably, for the object orientation, noise is first added to the Euler angles before converting them to quaternions which are finally used as the input to the policy. The details of each observation term, including noise ranges and scaling factors, are summarized in Table 5.
@@
Object State Position [0.01,0.01][-0.01,0.01] 1.0 mm
Linear Velocity [0.2,0.2][-0.2,0.2] 0.5 $m/s$
Orientation [0.05,0.05][-0.05,0.05] 1.0 radrad
Angular Velocity [0.2,0.2][-0.2,0.2] 0.25 $rad/s$
Robot Proprioception Projected Gravity Vector [0.05,0.05][-0.05,0.05] 1.0 -
Base Angular Velocity [0.2,0.2][-0.2,0.2] 0.25 $rad/s$
Joint Position [0.01,0.01][-0.01,0.01] 1.0 radrad
Joint Position [1.5,1.5][-1.5,1.5] 0.05 $rad/s$
@@
Last Action - 1.0 radrad
C.2 Curriculum of Velocity Commands
To encourage accurate and stable tracking, we adopt a curriculum that gradually expands the range of velocity commands as the policy achieves predefined thresholds in tracking performance and survival rate. At the beginning of training, both linear and angular velocity commands are constrained to small but non-zero ranges. These ranges are then symmetrically expanded step by step, with linear and angular components progressing independently until they reach their predefined maximum values. To maintain balanced development, we further restrict how far one curriculum can advance relative to the other: if the number of completed expansions for linear and angular ranges differs by more than a specified margin, further expansion is paused until the lagging dimension catches up.
C.3 Schedule of Zero Velocity Commands
As discussed in Section 5.2, we set the velocity commands to zero during the first $k$ steps of each episode ($k=50$ in our implementation) to ensure stable object placement and enable successful distillation from the teacher policy to the student policy with partial observations such as tactile sensing. This warm-up period allows the object to stabilize on the robot’s back before nonzero motion commands are applied. However, we observe that applying zero commands from the very beginning severely limits exploration, especially in the early phase of PPO training. Since most episodes terminate before reaching the $k$-th step, the policy is encouraged to maintain ground contact with all feet under zero commands, leading to premature convergence to relatively static behaviors rather than exploring long-swinging dynamic gaits.
To address this, we introduce a curriculum on the initial zero-command steps. At the beginning of training, we set the initial step length with zero commands to zero, which encourages a broader exploration of locomotion gaits. Once the command ranges are fully expanded and stabilized as described in Section C.2, we set this value to the final setting of $k=50$, ensuring that the teacher policy adapts well to start with balancing objects instead of directly transporting them. This scheduling strategy balances the need for early gait exploration with the later requirement of stable object transport. Alongside this schedule, we adjust the probability of sampling standing (zero-command) environments from a relatively higher value during early training to a smaller value in the final stage, ensuring sufficient locomotion diversity while still allowing a fraction of standing cases. Their values in our implementation are shown in Table 6.
Table 6: Curriculum schedule for zero velocity commands.
Parameter Value
Initial number of zero-command steps 0
Final number of zero-command steps $50$
Initial probability of sampling standing environments $0.10$
Final probability of sampling standing environments $0.05$
C.4 Reward Design
where $f_{i}$ is the $i$-th foot of the robot, $p^{w}_{i,z}$ is its height above the ground, $\mathbf{v}^{w}_{i,xy}$ is the lateral linear velocity of foot $i$ in the world frame, $h_{z,\text{th}}$ is the height threshold, and $v_{xy,\text{th}}$ is the lateral velocity threshold.